What is the current value of a $1,000 bond with a 6% annual coupon rate (paid semi-annually) that matures in 11 years if the appropriate stated annual discount rate is 12%.
Answer #1

Value of bond is equal to the present value of all future coupon payments and the principal amount

Semi annual discount rate = 12%/2 = 6%

Number of semi annual periods = 11*2 = 22 periods

Hence, value = 1000*6%*1/2*PVAF(6%, 22 periods) + 1,000*PVF(6%, 22 periods)

= 30*12.0416 +1,000*0.2775

=$638.748

i.e. $638.75